UT 20120601 Source Extractor Active Optics, AdSec Rested Collimation and Closed AO Loop on DX

Observer: DMiller, KBoutsia (LBTO) JCGuerra (Tucson)
Telescope Operator: DHuerta (LBTO)
Telescope Support: MMidkiff(LBTO)
Instrument Support: JMorris (LBTO)

Summary

Tasks:
  • Doug's IDL send binocular presets. It was quite easy.
  • Olga tested the binocular range balance. Trouble with authorize on left side.
  • Closed the DX AO loop on a bright star. This took about 2.5 hours as we had a hardware problem that required unplugging and plugging in the AGw/W Unit. Plus CCD39 stopped communicating, auto gain failed, re-rotator no enabled
  • Collect more Astig LUT and COW -vs- SEXT data

Details

02:15 Open chamber

Binocular presets from IDL

02:40 Try to send binocular preset. Received several warnings about requested instrument/focal station combination not allowed and binoculaControl SYNCPRESET command ignored: Duplicate event. Not clear to me the order of commands yet.

02:45 Success. Sent in acquire on both sides.

02:50 David align pointing model on left side. Could not see star on right side. Changing IE/CA on right. Left star stays stationary.

03:08 David re-initialize the right hexapod

03:15 Here's my syncpreset sequence
  • R iif_register
  • R iif_syncpreset
  • R iif_preset, ACT0277, MODE=ACQUIRE, GS=0
  • L iif_register
  • L iif_syncpreset
  • L iif_preset, ACT0277, MODE=ACQUIRE, GS=0

03:30 David send monocular preset to right side to find star

03:57 Find star and set IE/CA
  • Left: IE=-18.0, CA=-45.0
  • Right: IE = -54.5, CA = -137.8

03:58 Authorize on both sides. Send binocular preset. See star on right. Move star so can see on left.

04:05 Binocular preset to ACT0277, MODE=ACQUIRE, GS=1, PA=180. Success. A lot of coma on both sides. Collimating
  • Left: IE=-47.0, CA=-15.0
  • Right: IE = -49.5, CA = -147.8

Binocular Range Balance

04:15 Olga send several range bal commands. Getting error on left side about not registered

DX AO Closed Loop

04:53 Start with DX AO Closed Loop. Have several problems:
  • could not talk to ts8 in W Ring. Had to go to zenith, unplug and plug the power to the agw/W unit
  • lost communication to CCD37 twice. Had to restart slope computer
  • rerotator not getting rotator position. 10 mode closed loop cause shell to rip. Twice
  • autogain failed

07:00 Finally closed loop on AO1019. bin 1, 400 modes. offset, 5, 0 successful.

Going to collect LUT and Sext data

Astig LUT and Collimation data sets

New LUT and SEXT Collimation

Data in /OldRepository/AGW_Data/20120602/SEXTnewLUT

07:25 AO1018, el 79, wfsc Failed. David set pointing

07:31 AO1018, el 79, wfsc 1

07:36 AO883, el 42, wfsc 7

07:39 AO1029, el 67, wfsc 11

07:44 AO1039, el 29, wfsc 14

07:46 AO1044, el 49, wfsc 17

07:49 AO656, el 67, wfsc 22

07:53 AO696, el 28, wfsc 26

07:58 AO657, el 62, wfsc 31

08:03 AO733, el 45, wfsc 34

08:05 AO710, el 30, wfsc 37

08:09 AO731, el 54, wfsc 44

08:12 AO1028, el 77, wfsc 47

08:16 AO1041, el 40, wfsc 52

New LUT and COW Collimation

Data in /OldRepository/AGW_Data/20120602/COWnewLUT

08:24 AO1027, el 75

08:26 AO1041, el 45

08:29 AO657, el 68

08:33 AO711, el 32

08:36 AO735, el 50

08:39 AO657, el 70

08:42 AO780, el 28

08:47 AO729, el 64 saturated

08:51 AO731, el 62

08:53 AO1039, el 44

08:56 AO780, el 30

08:59 AO735, el 53

09:02 AO660, el 77

09:05 AO710, el 43

New LUT and SEXT Collimation, No Clear C00

Data in /OldRepository/AGW_Data/20120602/SEXTnewLUTnoC00

09:08 AO661, el 76, wfsc 1

09:11 AO711, el 41, wfsc 3

09:15 AO731, el 66, wfsc 7

09:18 AO1062, el 28, wfsc 10

09:22 AO1046, el 60, wfsc 16

09:26 AO1045, el 45, wfsc 20

09:30 AO803, el 30, wfsc 24

09:33 AO732, el 51, wfsc 27

09:35 AO661, el 81, wfsc 31

09:38 AO780, el 41, wfsc 34

09:42 AO1052, el 27, wfsc 37

09:44 AO1045, el 51, wfsc 39

09:47 AO731, el 72, wfsc 43

09:52 AO1064, el 40, wfsc 46

Old LUT and SEXT Collimation

Data in /OldRepository/AGW_Data/20120602/SEXToldLUT

09:56 AO731, el 74, wfsc 1

10:00 AO779, el 40, wfsc 4

10:05 AO735, el 66, wfsc 13

10:09 AO1052, el 33, wfsc 16

10:13 AO779, el 45, wfsc 22

10:16 AO735, el 69, wfsc 26

10:22 AO1069, el 29, wfsc 28

10:25 AO1039, el 60, wfsc 33

10:29 AO779, el 47, wfsc 37

10:32 AO763, el 30, wfsc 40

10:36 AO780, el 49, wfsc 46

10:39 AO733, el 75, wfsc 51

10:42 AO778, el 40, wfsc 53

New LUT and SEXT Image Quality

Data in /OldRepository/AGW_Data/20120602/IQ

10:49 AO735 clear C00, COW, el 77, collimate to 500, irtc 410, 110027

11:00 Clear C00, SEXT, el 77, collimate to 500, wfsc 1-2 wfsc camera problem

11:09 AO735, clear C00, SEXT, el 77, collimate to 500, wfsc still camera problems

11:25 David stopAGW and startAGW and now the camera works.

11:30 Getting too bright. Closing Chamber

-- DougMiller - 29 May 2012
Topic revision: r5 - 03 Jun 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback